Need help with this question. If you could can you please show your work. Thanks in advanced
Solve the linear system by Gauss-Jordan elimination.
y + z = 2
2x + 3z = 5
x + y + z = 3

Do you think this method has advantages over solving of simultaneous equations by earlier methods? Why or why not?

Answer by TimothyLamb(4379)

y + z = 2
2x + 3z = 5
x + y + z = 3
---
put the system of linear equations into standard form
---
y + z = 2
2x + 3z = 5
x + y + z = 3
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution by matrix method:
x = 1
y = 1
z = 1
---
now by elimination:
---
-3(y + z = 2)
2x + 3z = 5
---
-3y - 3z = -6
2x + 3z = 5
---
add the two equations:
---
-y = -1
y = 1
---
2x + 3z = 5
x + 1 + z = 3
---
2x + 3z = 5
-3(x + z = 2)
---
2x + 3z = 5
-3x - 3z = -6
---
add the two equations:
---
-x = -1
x = 1
---
y + z = 2
1 + z = 2
z = 1
---
solution by elimination:
x = 1
y = 1
z = 1
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---